# Rapid Gyroscope Calibration: A Deep Learning Approach

### v2

@@ V Conclusions
This study aimed to improve the zero-order calibration process of low-cost gyroscopes. To this end, we developed a simple yet efficient neural network for bias regression. To further improve performance, we incorporated MGs, using a mix of real and virtual data in the training dataset. We conducted the evaluation using 3 gyroscopes (a single IMU). For training, we created a dataset of real recorded data from 24 gyroscopes of two different brands and additional virtual data. The dataset contains 169 hours of recordings. The investigation revealed that our DL approach significantly reduced calibration time, i.e., the time to reach the performance achieved using the baseline approach, and significantly improved the accuracy of the model-based approach for the same calibration time. Given a single IMU, our DL calibration improved the baseline approach accuracy by $57\%$ and calibration time by $72\%$. With the addition of virtual data, improvement in accuracy increased further to $84\%$ and in calibration time to $79\%$. Using 12 MGs in the training set, greatly improved the results over a single IMU in both metrics, yet with 12 MGs, the addition of virtual data helped improve only the calibration time. The same behavior for 3 and 12 gyroscopes with and without virtual data was achieved using gyroscopes of a different brand, attesting to the robustness of our approach. Our findings reveal a critical trade-off between calibration speed and accuracy. With no time constraints, the model-based baseline approach continues to provide the most precise bias estimation. Yet, in applications such as search and rescue and robotics, where rapid calibration is a must, our approach has mission-critical advantages. Therefore, users can choose the most suitable method based on their concrete requirements, balancing the need for speed against accuracy and enhancing both flexibility and efficiency in various practical scenarios.

### v3

@@ V Conclusions
This study introduced a neural network-based approach to improve the zero-order calibration of low-cost gyroscopes by leveraging real and virtual data. Our deep learning method significantly reduced calibration time while improving accuracy compared to the baseline model-based approach. Specifically, using a single IMU, our approach achieved a 57% improvement in accuracy and a 72% reduction in calibration time. Incorporating virtual data further enhanced accuracy by 84% and reduced calibration time by 79%. When training with 12 MGs, the method showed substantial improvements in both metrics, with virtual data primarily contributing to faster calibration. These results remained consistent across different gyroscope brands, reinforcing the robustness of our approach.
Beyond its quantitative improvements, this method has significant practical implications. In applications where rapid calibration is crucial, such as robotics, autonomous vehicles, and search and rescue operations, our approach provides a viable alternative to traditional calibration methods, which require longer stationary periods. The ability to integrate virtual data also reduces the need for large scale real world sensor datasets, making this approach more scalable and adaptable to various hardware configurations. By bridging the gap between conventional calibration techniques and data-driven methods, our study contributes to the advancement of rapid and efficient inertial sensor calibration, with broad potential applications in navigation, robotics, and wearable technology.
However, our findings reveal a critical trade-off between calibration speed and accuracy. With no time constraints, the model-based baseline approach continues to provide the most precise bias estimation. Yet, in applications such as search and rescue and robotics, where rapid calibration is a must, our approach has critical mission advantages. Therefore, users can choose the most suitable method based on their concrete requirements, balancing the need for speed against accuracy and enhancing both flexibility and efficiency in various practical scenarios. Future work could explore extending this approach to dynamic calibration scenarios, improving real-time implementation, and optimizing the neural network for embedded systems to further enhance its applicability in resource-constrained environments.